OpenStudy (el_tucan):

Hi there. This is my question: Find an equation of the tangent line to the curve y = x * [x^(1/2)] that is parallel to the the line y = 1 + 3x. I'm able to find the derivative which I have as 1.5x^(1/2) but don't know where to go from there.

OpenStudy (el_tucan):

\[y = x \sqrt{x}\]

OpenStudy (el_tucan):

\[y' = (3/2) \sqrt{x}\]

OpenStudy (anonymous):

Yes, first you would need to find the DERIVATIVE of the curve. However you need to use the PRODUCT rule to find the derivative of \[x\]

OpenStudy (anonymous):

You need to use the PRODUCT rule to find the derivative of \[x \sqrt{x}\]

OpenStudy (anonymous):

do you know the product rule?

OpenStudy (campbell_st):

you don't need to product rule, by using index laws it can be simplified \[y = x \sqrt{x} = x^1 \times x^{\frac{1}{2}} = x^{\frac{3}{2}}\] so you have \[y = x^{\frac{3}{2}}\] so this makes finding the derivative so much easier

OpenStudy (el_tucan):

yes, i get \[y' = (3/2)\sqrt{x}\]

OpenStudy (anonymous):

@adziz is a smartass dont listen to him

OpenStudy (campbell_st):

ok... so your derivative is now equal to 3, which is the slope of the parallel line so \[3 = \frac{3}{2} \sqrt{x}\] find x...

OpenStudy (el_tucan):

ok, right!

OpenStudy (anonymous):

@adziz is a meanybutt

OpenStudy (campbell_st):

now when you have x, substitute it into the original equation to the find y and you have a point and a slope of 3 so you can find the equation of the tangent...

OpenStudy (anonymous):

@adziz

OpenStudy (anonymous):

ddd

OpenStudy (el_tucan):

sweet

OpenStudy (anonymous):

dsfsdffsfdhhfdhf

OpenStudy (el_tucan):

got it. thanks everyone.